Man is begins construction of a warehouse on July , 20X6. The project was completed on March 31, 20X7. No new loans were required to funds construction. Man is does have the following two interest-bearing liabilities that were outstanding throughout the sonctrcion period:

$2,000,00, 8% note

$8,000,000, 4% bonds

Construction expenditures incurred were as follow:

July 1, 20X6 $400,000

September 30, 20X6 $ 600,000

November 30, 20X8 $600,000

January 30, 20X7 $540,000

The Company’s Fiscal years-end is December 31.

Required:

Calcuate the amount of interest capitalized for 20X6 and 20X7.
Answer #1

1)Calculation of average interest rate:

Amount Rate Interest expense
Note 2000000 8% 2000000*8%=160000
Bond 8000000 4% 8000000*4%=320000
Total loan outstanding 10,000,000 Total interest 480000

Average interest rate =480000/10000000 = 4.80%

2)Calculation of weighted average capital expenditure for 2016:

Date Amount period of outstanding(out of 12 months) Average accumulated expenditure
1July 20x6 400000 6/12                 [1July-31Dec] 400000*6/12=200000

Sep 30 20x6

600000 3/12 600000*3/12=150000
Nov 30 20x6 600000 1/12 600000*1/12=50000
1600000
Total average expenditures 400000

Calculation of weighted average capital expenditure for 2016:

Date Amount period of outstanding(out of 12 months) Average accumulated expenditure
Beginning balance 1600000 3/12       [Jan -31March] 1600000*3/12=400000
30Jan 2017 540000 2/12      [Feb-31march] 90000
Total 490000

3)Calculation of interest capitalized :

for 2016 : Average accumulated expenditures*average rate

             = 400000*4.80%

             =19200

For 2017 : 490000*4.80%

              = 23520